Calculate the force of gravity between Earth and the moon
1
Expert's answer
2014-11-20T05:35:47-0500
Earth Mass = 5.9742 × 10^(24)
Moon Mass = 7.36 × 10^(22) kg
Distance Moon - Earth = 384,402 km
Universal Gravity Constant = 6.67*10^(-11)
F = Universal Gravity Constant * Earth Mass*Moon Mass /Distance Moon - Earth^2 = 1.985*10^26 N